# Leveraging Sparse Annotations for Leukemia Diagnosis on the Large Leukemia Dataset

### v1

@@ 5.1 Experiments on LeukemiaAttri
Implementation details: For AttriHead, dropout regularization (20%) is applied between FC layers, with Leaky ReLU activation used for both convolutional and FC layers. AttriDet is trained on the LeukemiaAttri dataset with specified augmentations and loss functions. We set the IoU threshold at 0.20 and a mosaic augmentation probability of 1.0. The optimizer settings include a weight decay of 0.0005, momentum of 0.937, and warmup biases of 0.1. Training is conducted for 100 epochs with a batch size of 4 on an NVIDIA GTX 1080 GPU, using a learning rate: of 0.01 to 0.001) and a 3-epoch warmup.
WBC Detection: To evaluate WBC detection performance on LeukemiaAttri, we perform experiments with several well-known object detectors including Sparse R-CNN [14], FCOS [15], DINO[16], and YOLOv5x [17]. We have performed detailed experiments on the dataset collected using low-cost microscope (LCM) and high-cost microscope (HCM) employing mobile cameras (C1) and special-purpose microscopic cameras (HD1500T for HCM and MZZCAT 5MP for LCM - C2) on 40x and 100x magnifications. The experimental results are shown in Table 4. The results indicate that the best WBC detection accuracy (mAP@50-95 of 26.3 and mAP@50 of 44.2) is achieved on the data collected using HCM at 100x employing C2 camera using YOLOv5x. Similarly, when comparing on 40x, YOLOv5x outperforms ($\text{mAP}_{50\text{-}95}$ of 20.1 and $\text{mAP}_{50}$ of 37.3) other objection detection methods. A similar pattern can be observed for other subsets as well. We believe this is due to the YOLOV5 robust feature learning for different sizes of cells in the LeukemiaAttri dataset. Given that YOLOv5x demonstrated superior performance across both the HCM and LCM subsets of microscope and mobile camera data, we extend this for attribute prediction.
Table 4: Object Detection baselines results ($mAP_{@50}$) on multipal subset of LeukemiaAttri dataset. The C2 camera represents the HD1500T camera for HCM, while the C2* camera corresponds to the ZZCAT 5MP camera for LCM. The C1 mobile camera, used for both, is Honor 9x Lite.
@@
Sparse R-CNN [14] 29.6 36.7 27.0 32.7 32.6 25.9 26.4 33.9
FCOS [15] 31.8 40.6 24.8 32.7 33.9 34.3 31.2 28.5
DINO[16] 33.8 43.7 36.1 36.9 34.3 38.2 31.4 36.6
YOLOv5x[17] 38.8 44.2 36.1 37.3 39.5 29.7 38.1 34.9
WBC detection with attribute prediction: Table. 5 demonstrate results of our proposed AttriDet for attributes prediction. In addition, we demonstrate the improved WBC detection of AttriDet (last column) as compared to standard YOLOv5. We credit the enhanced WBC results to robust feature learning using the attributes head. Findings from CBM [48] and AttriDet demonstrate superior performance in predicting WBC types and their attributes. However, detecting nucleus and cytoplasmic vacuoles remains a significant challenge.
Table 5: Testing set results of AttriDet and SOTA methods on HCM_100x_C2 subset: WBC Type, Attributes (NC: Nuclear Chromatin, NS: Nuclear shape, N: Nucleus, C: Cytoplasm CB: Cytoplasmic basophilia CV: Cytoplasmic vacuoles)
Method F1 WBC
NC NS N C CB CV mAP@50-95
CBM [48] 21.9 96.2 41.8 77.2 70.2 3.33 27.6
AttriDet [19] 73.9 95.9 54.3 89.7 83.6 29.1 28.2
Unsupervised Domain Adaptation based WBC detection: The LeukemiaAttri dataset includes 12 subsets originating from distinct domains, captured via HCM and LCM. These subsets exhibit significant domain shifts, as illustrated in Fig. 2 (a), positioning our dataset as a valuable benchmark for domain adaptation studies. It is worth noting that LCM images often suffer from poor quality, discouraging hematologists from providing annotations due to the labor-intensive and error-prone nature of the process. As a solution, UDA approaches enable the training of object detectors on high-quality, well-annotated HCM images and facilitate their application to LCM images.
@@
      YOLOv5[17]       11.0       25.5
      DACA [49]       12.6       30.2
      ConfMix [14]       12.6       33.5
To establish the UDA baselines, we evaluated UDA baselines using ConfMix [14] and DACA [49] on high-resolution subsets (HCM_100x_C2 as source and LCM_100x_C2 as target). Table 6 demonstrates the YOLOv5x (source only) achieved 25.5 mAP@50, while UDA methods improved performance to 33.5 (ConfMix) and 30.2 (DACA), highlighting the significant domain shift and complexity of our dataset.
5.2 Experiments on Sparse-LeukemiaAttri
Implementation details: For the SLA-Det implementation, we set the values as follows: $\mathbf{t}_{0}:0.70$, $\mathbf{t}_{1}:0.95$, $\mathbf{t}_{2}$: 0.26. Additionally, the batch size is set to 4, and both pseudo and triplet losses are used with a gain of 0.1. During training for WBC detection, the triplet loss is utilized for 40 epochs, while the other loss functions are applied for the entire 50 epochs. In the case of detection with attribute prediction, the triplet loss is activated for 10 epochs, while the others for 30.
WBC Detection: We compare the proposed SLA-Det with recent state-of-the-art sparse learning methods [33, 36]. The results in Table 7 demonstrate that our approach significantly outperforms existing methods, achieving notably higher mAP. The superior results of our approach are due to effectively utilizing annotated information through labeled region masking, strategic use of pseudo-labels, and integrating cell similarity.
@@ 5.2 Experiments on Sparse-LeukemiaAttri
Ablation Studies: Our method contains several components. We analyze the components of the SLA-Det approach for WBC detection to verify their effectiveness. Results in Table 8 show that each component contributes to the final accuracy, with LR loss improving detection mAP by 3.1%. PL loss achieves an additional 0.7% improvement and applying triplet loss, results in a significant 3.6% increase. Our results demonstrate that each component of our approach contributes effectively.
Table 8: Ablation results of SLA-Det on Sparse-LeukemiaAttri
Labeled-Region Pseudo-label Triplet mAP@50-95 mAP@50
24.9 40.8
25.1 41.5
27.2 45.1
We have also tried different straightforward methods to address the sparse dataset problem. In the first experiment, the model is trained on cropped, annotated patches from Sparse-LeukemiaAttri, while during testing, each image is divided into overlapping patches. In the second experiment, the model is trained on the full image with unannotated regions masked out, and the entire image is used during testing. Both experiments yielded $mAP@{{}_{50}}$ scores of 37.7% and 37.6%, respectively, which are noticeably lower than the 45.1% mAP@50 achieved by our proposed approach.
Table 9: Ablation study results: F1 Score of SLA-Det on Sparse-LeukemiaAttri. Attributes are the same as in Table 5
@@
SLA-Det* 71.2 98.0 50.9 87.8 84.4 14.0 27.6
SLA-Det** 71.3 97.7 51.3 87.7 84.3 10.8 28.1
SLA-Det*** 70.9 97.7 51.3 87.9 84.3 12.1 29.0
@@
WBC detection with attribute prediction : Table. 9 presents the results of our SLA-Det model for various prediction of attributes. In this table, SLA-Det* denotes the AttriHead with only the LR loss. SLA-Det** includes both LR and PL losses with the AttriHead, representing the second phase of the ablation study. Finally, SLA-Det*** shows the results from the final phase of the ablation study, incorporating all losses (LR+PL+Triplet) with the AttriHead.
@@
Figure 6: Qualitative comparison of different sparsely annotated approaches, including SparseDet [33] and Co-mining [36], and SLA-Det. Previous methods struggle to accurately localize all WBCs (columns 2 and 4) whereas the proposed approach effectively detects WBCs with improved classification. A zoomed-in view highlights (a) successful detection cases and (b) misclassified cases labeled as None (details are given in text).
@@
Qualitative Results: The qualitative comparison of the proposed approach and existing sparse object detectors [36, 33] is shown in Figure 6. The results show that the proposed method’s localization and classification capability significantly outperforms the competitive approaches. Figure 6 (a) and Figure 6 (b) show the zoomed-in version of some of the cells. In Figure 6 (a) all three cells belong to the same myeloid lineage and exhibit nearly identical morphological features. The primary difference between myeloblasts and monoblasts is the nucleus’s position within the cytoplasm: monoblasts have a nucleus surrounded by a more abundant and vacuolated cytoplasm, while myeloblasts have a more distinct nuclear-cytoplasmic boundary. Despite these minor differences, the detector effectively identifies both cell types Figure 6 (b) demonstrates a failure case of our approach, where the shapes of the cell nucleus and cytoplasm are highly irregular. The first two cells are labeled as ‘Monocyte’ and the third cell is labeled as ‘None’. We believe that the minor staining effects might be the reason to confuse the detector, potentially resulting in incorrect classifications.
@@ 6 Conclusion
We present a Large Leukemia Dataset (LLD) consisting of datasets named LeukemiaAttri  and Sparse-LeukemiaAttri. The LeukemiaAttri  contains 12 subsets collected by using two different quality microscopes with multiple cameras at different resolutions (10x, 40x, and 100x). There are 2.4k annotated images per resolution (in both HCM and LCM), with 10.3K cells annotated for 14 WBC types and 7 morphological properties. Based on morphological information, we have provided an AttriDet method to detect the WBC type with its morphological attributes. The Sparse-LeukemiaAttri is introduced to tackle the real-world challenges of sparse annotation in microscopy images. To establish the baseline solutions, we propose SLA-Det, which effectively utilizes labeled and unlabeled region information. We believe that the presented dataset and methods will advance research in explainable, robust, and generalized leukemia detection.

### v2

@@ 5.1 Experiments on LeukemiaAttri
Implementation details: For AttriHead, dropout regularization (20%) is applied between FC layers, with Leaky ReLU activation used for both convolutional and FC layers. AttriDet is trained on the LeukemiaAttri dataset with specified augmentations and loss functions. We set the IoU threshold at 0.20 and a mosaic augmentation probability of 1.0. The optimizer settings include a weight decay of 0.0005, momentum of 0.937, and warmup biases of 0.1. Training is conducted for 100 epochs with a batch size of 4 on an NVIDIA GTX 1080 GPU, using a learning rate: of (0.01 to 0.001) and a 3-epoch warmup.
WBC detection: To evaluate WBC detection performance on LeukemiaAttri, we perform experiments with several well-known object detectors including Sparse R-CNN Sun et al. (2021), FCOS Tian et al. (2019), DINOZhang et al. (2022), and YOLOv5x Ultralytics (2021). We have performed detailed experiments on the dataset collected using a low-cost microscope (LCM) and a high-cost microscope (HCM), employing mobile cameras (C1) and special-purpose microscopic cameras (HD1500T for HCM and MZZCAT 5MP for LCM - C2) on 40x and 100x magnifications. The experimental results are shown in Table 4. The results indicate that the best WBC detection accuracy (mAP@50-95 of 26.3 and mAP@50 of 44.2) is achieved on the data collected using HCM at 100x employing C2 camera using YOLOv5x. Similarly, when comparing on 40x, YOLOv5x outperforms ($\text{mAP}_{50\text{-}95}$ of 20.1 and $\text{mAP}_{50}$ of 37.3) other objection detection methods. A similar pattern can be observed for other subsets as well. We believe this is due to the YOLOV5 robust feature learning for different sizes of cells in the LeukemiaAttri dataset. Given that YOLOv5x demonstrated superior performance across both the HCM and LCM subsets of microscope and mobile camera data, we extend this for attribute prediction.
Table 4: Object Detection baselines results ($mAP_{@50}$) on multipal subset of LeukemiaAttri dataset. The C2 camera represents the HD1500T camera for HCM, while the C2* camera corresponds to the ZZCAT 5MP camera for LCM. The C1 mobile camera, used for both, is the Honor 9x Lite.
@@
Sparse R-CNN 29.6 36.7 27.0 32.7 32.6 25.9 26.4 33.9
FCOS 31.8 40.6 24.8 32.7 33.9 34.3 31.2 28.5
DINO 33.8 43.7 36.1 36.9 34.3 38.2 31.4 36.6
YOLOv5x 38.8 44.2 36.1 37.3 39.5 38.1 29.7 34.9
WBC detection with attribute prediction: Table. 5 demonstrate results of our proposed AttriDet for attributes prediction. In addition, we demonstrate the improved WBC detection of AttriDet (last column) as compared to standard YOLOv5. We credit the enhanced WBC results to robust feature learning using the attributes head. Findings from CBM Keser et al. (2023) and AttriDet demonstrate superior performance in predicting WBC types and their attributes. However, detecting nucleus and cytoplasmic vacuoles remains a significant challenge.
[…]
Method NC NS N C CB CV WBC
CBM 21.9 96.2 41.8 77.2 70.2 3.3 27.6
AttriDet 73.9 95.9 54.3 89.7 83.6 29.1 28.2
Ablation of attrihead. The ablation of the AttriHead architecture could offer useful insights. Our focus was on developing a lightweight model tailored to real-world clinical constraints and limited computational resources. The chosen design, guided by empirical observations, balances performance and efficiency, as demonstrated by the strong results reported in Table 6 under low resource settings.
Table 6: Ablation study of activation functions and architectural configurations of Attrihead on attribute-level F1 scores.
Activation FC Conv NC NS N C CB CV
Layers Layers
LeakyReLU 3 3 73.9 95.9 54.3 89.7 83.6 29.1
ReLU 3 3 64.1 96.2 47.8 88.1 83.8 11.2
LeakyReLU 3 2 65.8 96.3 42.8 87.7 84.2 04.7
LeakyReLU 2 3 63.8 96.0 50.2 89.7 83.5 2.3
LeakyReLU 2 2 65.0 96.3 45.6 88.1 84.3 10.8
Unsupervised domain adaptation based WBC detection:
The LeukemiaAttri dataset includes 12 subsets originating from distinct domains, captured via HCM and LCM. These subsets exhibit significant domain shifts, as illustrated in Fig. 2 (a), positioning our dataset as a valuable benchmark for domain adaptation studies. It is worth noting that LCM images often suffer from poor quality, discouraging hematologists from providing annotations due to the labor-intensive and error-prone nature of the process. As a solution, UDA approaches enable the training of object detectors on high-quality, well-annotated HCM images and facilitate their application to LCM images.
@@
YOLOv5 11.0 25.5
DACA 12.6 30.2
ConfMix 12.6 33.5
To establish the UDA baselines, we evaluated UDA baselines using ConfMix Sun et al. (2021) and DACA Mekhalfi et al. (2023) on high-resolution subsets (HCM_100x_C2 as source and LCM_100x_C2 as target). Table 7 demonstrates that YOLOv5x (source only) achieved 25.5 mAP@50, while UDA methods improved performance to 33.5 (ConfMix) and 30.2 (DACA), highlighting the significant domain shift and complexity of our dataset.
@@
5.2 Experiments on Sparse-LeukemiaAttri
Implementation details: For the SLA-Det implementation, we set the values as follows: $\mathbf{t}_{0}:0.70$, $\mathbf{t}_{1}:0.95$, $\mathbf{t}_{2}$: 2.6. Additionally, the batch size is set to 4, and both pseudo and triplet losses are used with a weight of 0.1. During training for WBC detection, the triplet loss is utilized for 40 epochs, while the other loss functions are applied for the entire 50 epochs. In the case of detection with attribute prediction, the triplet loss is activated for 10 epochs, while the others are for 30. In all of the experiments, we used the final model weights (i.e., the weights from the last training epoch) for evaluation. The same dataset split and annotations are used for our approach and the baseline methods.
In our setting, approximately 20% regions per image are annotated, and the rest remain unannotated. For the SOTA methods, we strictly followed their official implementation and default hyperparameters as provided in their publicly available repositories.
WBC detection: We compare the proposed SLA-Det with recent state-of-the-art sparse learning methods Suri et al. (2023); Wang et al. (2021). The results in Table 8 demonstrate that our approach outperforms existing methods, achieving a notably higher mAP. The superior results of our approach are due to effectively utilizing annotated information through labeled region masking, strategic use of pseudo-labels, and integrating cell similarity.
Table 9: Ablation results of SLA-Det on Sparse-LeukemiaAttrifor detection only, we conduct multiple runs (N=3) for each ablation setting and compute the mean and standard deviation of both mAP@50-95 and mAP@50.
Labeled Pseudo Triplet mAP@50-95 mAP@50
Region label (±std) (±std)
$24.6_{(\pm 0.5)}$ $40.6_{(\pm 0.4)}$
$24.8_{(\pm 0.2)}$ $41.3_{(\pm 0.4)}$
$\mathbf{26.4}_{(\pm 0.4)}$ $\mathbf{44.6}_{(\pm 0.5)}$
Ablation studies: Our method contains several components. We analyze the components of the SLA-Det approach for WBC detection to verify their effectiveness. Results in Table 9 show that each component contributes to the final accuracy, with LR loss improving detection mAP by 3.1%. PL loss achieves an additional 0.7% improvement, and applying triplet loss results in a significant 3.6% increase. Our results demonstrate that each component of our approach contributes effectively.
Table 10: Ablation study results: F1 Score of SLA-Det on Sparse-LeukemiaAttri. Attributes are the same as in Table 5. Multiple runs (N=3) are conducted for each ablation setting and compute the mean and standard deviation of mAP@50-95.
@@ 5.2 Experiments on Sparse-LeukemiaAttri
SLA-Det* $70.8_{(\pm 0.6)}$ $95.5_{(\pm 0.07)}$ $50.9_{(\pm 0.07)}$ $88.3_{(\pm 0.7)}$ $83.9_{(\pm 0.6)}$ $14.7_{(\pm 0.9)}$ $26.4_{(\pm 0.5)}$
SLA-Det** $69.1_{(\pm 1.9)}$ $97.5_{(\pm 0.2)}$ $50.4_{(\pm 1.1)}$ $87.8_{(\pm 0.7)}$ $83.6_{(\pm 0.7)}$ $11.9_{(\pm 3.1)}$ $26.9_{(\pm 0.7)}$
SLA-Det*** $\textbf{71.1}_{(\pm 1.2)}$ $\textbf{97.5}_{(\pm 0.1)}$ $\textbf{51.2}_{(\pm 0.8)}$ $\textbf{88.0}_{(\pm 0.3)}$ $\textbf{84.2}_{(\pm 0.2)}$ $\textbf{13.2}_{(\pm 1)}$ $\textbf{28.2}_{(\pm 0.7)}$
We have also tried different straightforward methods to address the sparse dataset problem. In the first experiment, the model is trained on cropped, annotated patches from Sparse-LeukemiaAttri, while during testing, each image is divided into overlapping patches. In the second experiment, the model is trained on the full image with unannotated regions masked out, and the entire image is used during testing. Both experiments yielded $mAP@{{}_{50}}$ scores of 37.7% and 37.6%, respectively, which are noticeably lower than the 45.1% mAP@50 maximum achieved by our proposed approach.
@@
WBC detection with attribute prediction : Table 10 presents the results of our SLA-Det model for various prediction of attributes. In this table, SLA-Det* denotes the AttriHead with only the LR loss. SLA-Det** includes both LR and PL losses with the AttriHead, representing the second phase of the ablation study. Finally, SLA-Det*** shows the results from the final phase of the ablation study, incorporating all losses (LR+PL+Triplet) with the AttriHead.
Ablation of thresholds : The thresholds $t_{0}$, $t_{1}$, and $t_{2}$ used in the pseudo-label filtering step of SLA-Det play a critical role in influencing detection performance. These values were selected empirically based on preliminary experiments. To quantify their impact, an ablation was conducted by varying each threshold within a reasonable range. The results, presented in Table 11, demonstrate that our method is robust to slight changes in these parameters.
Table 11: Ablation of pseudo-label thresholds ($t_{0}$, $t_{1}$, $t_{2}$) on SLA-Det performance.
$t_{0}$ $t_{1}$ $t_{2}$ NC NS N C CB CV $mAP@_{50-95}$ $mAP@_{50}$
0.70 0.95 2.6 70.9 97.7 51.3 87.9 84.3 12.1 29.0 46.4
0.70 0.93 2.6 68.4 97.6 49.9 89.5 83.3 13.2 25.5 41.5
0.70 0.90 2.6 70.9 97.7 52.5 88.6 83.4 11.6 19.6 33.2
0.60 0.95 2.6 71.6 97.3 52.1 89.5 84.8 13.1 26.9 44.3
0.70 0.95 2.5 70.8 97.1 51.7 89.1 83.1 10.3 25.7 42.4
0.70 0.95 2.7 70.3 97.6 50.5 89.0 84.0 8.0 26.4 42.9
Discussion: The model’s difficulty in accurately detecting the ‘Cytoplasmic Vacuoles’ attribute is primarily due to two main factors. First, there is a significant class imbalance in the dataset (see Table 3). Second, cytoplasmic vacuoles are morphologically linked to cell swelling and early necrotic changes, which are often difficult to distinguish under variable staining and imaging conditions Chiang et al. (2018). Their nonspecific appearance across multiple pathological states and overlap with other intracellular features further adds to annotation and detection challenges. In routine clinical workflows, clinicians typically perform a quick differential count in a few minutes; however, detailed morphological assessment requires expertise and is time-consuming. A model like SLA-Det, trained on a dataset with annotated morphological attributes, can assist hematologists in real time during inference by providing cell type and morphology insights as a second opinion. While only $\approx 20\%$ of each image is annotated for training to reduce annotation burden, the model is designed to operate on the entire field of view (FoV) during inference. With an average processing time of $\approx 0.29$ seconds per image ($\approx 3.45$ frames per second), the model operates at near real-time performance suitable for clinical workflows.
@@
Figure 6: Qualitative comparison of different sparsely annotated approaches, including SparseDet Suri et al. (2023) and Co-mining Wang et al. (2021), and SLA-Det. Previous methods struggle to accurately localize all WBCs (columns 2 and 4), whereas the proposed approach effectively detects WBCs with improved classification. A zoomed-in view highlights (a) successful detection cases and (b) misclassified cases labeled as None (details are given in the text).
@@
Qualitative results: The qualitative comparison of the proposed approach and existing sparse object detectors Wang et al. (2021); Suri et al. (2023) is shown in Figure 6. The results suggest that the proposed method offers improved localization and classification performance compared to competitive approaches. Figure 6 (a) and Figure 6 (b) show the zoomed-in version of some of the cells. In Figure 6 (a), all three cells belong to the same myeloid lineage and exhibit nearly identical morphological features. The primary difference between myeloblasts and monoblasts is the nucleus’s position within the cytoplasm: monoblasts have a nucleus surrounded by a more abundant and vacuolated cytoplasm, while myeloblasts have a more distinct nuclear-cytoplasmic boundary. Despite these minor differences, the detector effectively identifies both cell types. Figure 6 (b) demonstrates a failure case of our approach, where the shapes of the cell nucleus and cytoplasm are highly irregular. The first two cells are labeled as ‘Monocyte’ and the third cell is labeled as ‘None’. We believe that the minor staining effects might be the reason for confusing the detector, potentially resulting in incorrect classifications.
@@ 6 Conclusion
We present a Large Leukemia Dataset (LLD) consisting of datasets named LeukemiaAttri  and Sparse-LeukemiaAttri. The LeukemiaAttri  contains 12 subsets collected by using two different quality microscopes with multiple cameras at different resolutions (10x, 40x, and 100x). There are 2.4k annotated images per resolution (in both HCM and LCM), with 10.3K cells annotated for 14 WBC types and 7 morphological properties. Based on morphological information, we have provided an AttriDet method to detect the WBC type with its morphological attributes. The Sparse-LeukemiaAttri is introduced to tackle the real-world challenges of sparse annotation in microscopy images. To establish the baseline solutions, we propose SLA-Det, which effectively utilizes labeled and unlabeled region information. We believe that the presented dataset and methods will advance research in explainable, robust, and generalizable leukemia detection. Our propose setup ensures a clean and consistent sparse annotation structure. Real-world clinical annotations are often irregular and unstructured, with cells marked at arbitrary locations. The current SLA-Det model does not support unstructured annotation patterns. Addressing this limitation will be a focus of future work.